# Argumentation Computation with Large Language Models : A Benchmark Study

### v1

@@ Abstract
In recent years, large language models (LLMs) have made significant advancements in neuro-symbolic computing. However, the combination of LLM with argumentation computation remains an underexplored domain, despite its considerable potential for real-world applications requiring defeasible reasoning. In this paper, we aim to investigate the capability of LLMs in determining the extensions of various abstract argumentation semantics. To achieve this, we develop and curate a benchmark comprising diverse abstract argumentation frameworks, accompanied by detailed explanations of algorithms for computing extensions. Subsequently, we fine-tune LLMs on the proposed benchmark, focusing on two fundamental extension-solving tasks. As a comparative baseline, LLMs are evaluated using a chain-of-thought approach, where they struggle to accurately compute semantics. In the experiments, we demonstrate that the process explanation plays a crucial role in semantics computation learning. Models trained with explanations show superior generalization accuracy compared to those trained solely with question-answer pairs. Furthermore, by leveraging the self-explanation capabilities of LLMs, our approach provides detailed illustrations that mitigate the lack of transparency typically associated with neural networks. Our findings contribute to the broader understanding of LLMs’ potential in argumentation computation, offering promising avenues for further research in this domain.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have significantly improved their proficiency in addressing a range of symbolic computation tasks and complex reasoning problems, thereby demonstrating their adaptability across diverse domains Fang et al. (2024); Huang et al. (2022); Alon et al. (2022); Pan et al. (2023). Nevertheless, LLMs exhibit comparatively limited efficacy in formal logical reasoning tasks Han et al. (2022); Eisape et al. (2024), frequently resulting in hallucinations and derivation errors. Integrating LLMs with symbolic methods remains a pivotal challenge in the advancement of contemporary artificial intelligence systems.
Argumentation is a form of defeasible reasoning that plays a crucial role in various real-world applications, such as legal decision-making Prakken and Sartor (2023), policy compliance analysis Dong et al. (2022), and ethical decision-making Awad et al. (2020). Combining argumentation with deep learning and gaining insights into how humans engage in argumentative reasoning is a vital aspect of artificial intelligence development. Previous research has pioneered this area by employing neural networks to compute acceptability semantics within the abstract argumentation framework (AAF) Dung (1995). Notable contributions include AGNN and EGNN  Craandijk and Bex (2020, 2022), which explores the use of graph neural network (GNN) to determine argumentation semantics, demonstrating that neural networks can effectively learn and predict the acceptability of arguments in various scenarios. These works highlight the promising intersection of neural networks and argumentation theory, however, the intrinsic black-box characteristic of neural networks leads to the lack of explainability in prediction. It is hard to show explicitly how their models predict the acceptability as their basis for decision-making, which is crucial in real applications.
In this work, we focus on exploring computational problems within the abstract argumentation framework, with an emphasis on leveraging LLMs to advance research in this domain. Our objectives are twofold: first, to train an LLM capable of both computing semantics and explaining the underlying algorithm process; second, to enhance the explainability of LLMs through the training, potentially overcoming some of their inherent limitations. To achieve these goals, we utilize the self-explanation capacity of LLMs to fine-tune language models on a curated data corpus. This approach enables LLM to interpret its behavior during the problem-solving process through text generation. Specifically, we establish an explainable AAF computation benchmark containing frameworks of varying difficulty levels. We explore two fundamental tasks: computing the grounded extensions and complete extensions based on labelling algorithms. In the dataset, each sample comprises an AAF described using a graph description language, with the execution information of a recursive algorithm and justification process attached as the explanation. The data organization embeds a post-explainability mechanism into the training process, which can reflect the LLM’s capability in understanding symbolic computation.
To summarize, our contributions are as follows:
  • We establish an explainable argumentation computation benchmark, enabling the training of explainable LLMs to compute extensions using labeling algorithms.
  • A methodology for integrating argumentation frameworks with language models is proposed through fine-tuning, improving the model’s ability to compute semantics and explain underlying computation dynamics.
2 Preliminaries
@@
Abstract argumentation frameworks can be concisely represented by directed graphs, where arguments are nodes and edges represent the attack relation. We use the AAF in Figure 1 (denoted as Example 1) as our illustration example in this paper. The status of a given argument is determined through argumentation semantics, often yielding results in the form of extensions Dung (1995) or labellings Wu et al. (2010), which can be used interchangeably. In this work, we primarily use labellings, and the explanation of computation also involves concepts such as conflict-freeness.
@@
Definition 2.
A labelling $\mathcal{L}$ for an abstract argumentation framework $\mathcal{F}=(\mathcal{A},\mathcal{R})$ is a function $\mathcal{L}:\mathcal{A}\rightarrow\{\mathtt{IN},\mathtt{OUT},\mathtt{UNDEC}\}$.
@@ 3 Dataset Construction
To construct a benchmark with algorithm explanations, we first generate randomly a variety of challenging argumentation frameworks following the methodology described in Craandijk and Bex (2020). In generation process, duplicate samples are removed and the ground truth semantics are computed. The number of arguments in an AAF ranges from 6 to 25. To encompass various levels of difficulty, we generate 3,000 samples for the training set and 100 samples for the test set for each number of arguments $n$, resulting in a total of 60,000 training samples and 2,000 test samples. The grounded, preferred, stable, and complete semantics have an average of 1.0, 1.4, 1.1, and 2.4 extensions per AAF, respectively, with 4.3, 6.4, 7.6, and 6.0 arguments per extension, respectively.
@@ 3.1 Instruction and Problem
The instruction includes a basic introduction of AAF, the task description, and return format requirements. Since we are conducting formal reasoning, we ask LLMs to generate a JSON object at the end for clarity and concreteness. To render the abstract representation readable for both humans and LLMs, we adopt graph description languages to represent the original argumentation frameworks. We use three graph description languages: GraphML Brandes et al. (2002), Graphviz DOT, and JSON. For each data sample, one format is randomly chosen during the problem statement generation.
@@ 3.2.1 Grounded Semantics
The labelling algorithm for generating the grounded labelling begins by assigning $\mathtt{IN}$ to all arguments that are not attacked. It then proceeds iteratively: $\mathtt{OUT}$ is assigned to any argument attacked by an argument that has just been made $\mathtt{IN}$, and $\mathtt{IN}$ is assigned to arguments whose attackers are all $\mathtt{OUT}$. The iteration continues until no new arguments can be assigned $\mathtt{IN}$ or $\mathtt{OUT}$. Any arguments that remain unlabelled are then assigned $\mathtt{UNDEC}$. This algorithm is sound and complete as it effectively mimics the construction of the grounded extension through the iteration of a framework’s characteristic function Modgil and Caminada (2009).
@@ 3.4 Enhancing Data Diversity
To increase the data diversity and mitigate overfitting, we employ GPT-4o OpenAI (2023) to re-illustrate the entire text for both training and testing samples. As a necessary requirement, we maintain the formal part in each sample to ensure the validity and consistency of the data.
4 Chain-of-Thought Baseline
Prompting generally results in models making predictions with lower accuracy compared to models that have been fine-tuned with substantial training data Zhang et al. (2023). To validate this observation and establish a baseline method, we design a Chain-of-Thought (CoT) pipeline Wei et al. (2022) that computes extensions based on a labelling algorithm. The argumentation computation is divided into multiple steps, framed as a reasoning task.
For computing grounded semantics, LLMs are taught to identify arguments and their relations, starting with those that aren’t attacked. They follow steps to label arguments until no more changes are needed. As for complete Semantics, the models take arguments that are labelled $\mathtt{UNDEC}$ and decide which can be further add to set $\mathtt{IN}$ to form a complete extension, checking if their choices make sense. The concrete prompts and implementation steps can be found in Appendix in supplementary materials.
@@ 4 Chain-of-Thought Baseline
Qwen2-Train 0.91 0.98 0.79 0.99 0.87 0.80 0.88 0.87 0.83 0.85 0.85
Llama-3-Train 0.95 0.99 0.85 0.99 0.89 0.86 0.90 0.89 0.88 0.88 0.88
@@
5 Experiment
Figure 4: Learning curves.
5.1 Implementation Details
In our training process, we employ the LoRA technique Hu et al. (2022) to facilitate efficient fine-tuning. Throughout all experiments, we utilize the bfloat16 (bf16) computation datatype. The learning rate is set to $410e-$ 4 and the maximum context length to 2048 tokens. The LLMs are trained for 8 epochs using a server with eight RTX 4090 24GB GPUs, which required approximately 20 hours.
To assess the performance of models, we consider two perspectives: the view of extension and the view of argument. The accuracy of extension prediction, denoted as ACCσ, is the proportion of samples where the predicted $\mathtt{IN}(\mathcal{L})$ exactly matches the ground truth extension. During inference, we calculate argument credulous/skeptical acceptability based on predicted labellings. For argument acceptability classification, we use the Matthews Correlation Coefficient (MCC) as the metric due to the imbalance between positive and negative samples, denoted as MCCc and MCCs respectively.
@@ 5.2 Main Results
It is important to acknowledge that the accuracy of our models is comparatively lower than that achieved by GNN-based methods Craandijk and Bex (2020) which use trained multiple models for the prediction of accepted arguments under different semantics, since GNNs are inherently more suited to argumentation computation due to their architectural design. However, a key advantage of our method lies in its ability to directly compute extensions and provide corresponding explanations, offering a comprehensive understanding of the argumentation process. This capability underscores the practical applicability of our model in diverse semantic contexts.

### v2

# Enhancing Conflict Resolution in Language Models via Abstract Argumentation

@@ Abstract
In recent years, large language models (LLMs) have made significant advancements in developing human-like and engaging dialogue systems. However, in tasks such as consensus-building and persuasion, LLMs often struggle to resolve conflicts arising from incomplete or inconsistent information, revealing their limitations in real-world applications. Given these limitations, abstract argumentation, a specialized logical framework designed to resolve conflicts and inconsistencies, becomes particularly relevant. In this paper, we aim to enhance the conflict-solving capabilities of LLMs by leveraging formal abstract argumentation, integrating language model learning with symbolic computation. To achieve this, we develop and curate a dataset comprising diverse abstract argumentation frameworks, accompanied by detailed explanations of the argument acceptability computation process. Subsequently, we fine-tune LLMs on this dataset, focusing on abstract conflict resolution tasks. As a comparative baseline, LLMs are also evaluated using a chain-of-thought approach, however, they fail to solve the conflict-based arguments effectively. Our experiments demonstrate that process explanations play a crucial role in learning. Models trained with explanations exhibit superior generalization accuracy compared to those trained solely on question-answer pairs. Furthermore, leveraging LLMs’ self-explanation capabilities, our approach provides detailed illustrations that mitigate the lack of transparency typically associated with neural networks.
journal: Neurocomputing
[…]
@@ 1 Introduction
Recent advancements in large language models (LLMs) have significantly improved their proficiency in addressing a range of complex reasoning problems, thereby demonstrating their adaptability across diverse domains [fang2024large, huang2022language, alon2022neuro, pan-etal-2023-logic]. Nevertheless, LLMs exhibit comparatively limited efficacy in formal logical reasoning tasks [han2022folio, li2025exploring, eisape2024systematic, cheng2025empowering], frequently resulting in hallucinations and derivation errors. These characteristics render them unreliable in applications such as conflict resolution and mediation. Therefore, it is desirable to enhance LLMs’ ability to perform reasoning and make relatively sound decisions when dealing with complex and conflicting information.
Argumentation is a formal form of defeasible reasoning that plays a crucial role in various real-world applications, such as legal decision-making [prakken2023formal, chi2021optimized], policy compliance analysis [dong2022defeasible], and ethical decision-making [awad2020approach]. Combining argumentation with deep learning and gaining insights into how humans engage in argumentative reasoning is a vital aspect of artificial intelligence development. Previous research has pioneered this area by employing neural networks to compute acceptability semantics within the abstract argumentation framework (AAF) [dung1995acceptability, malmqvist2024approximating, Thorburn2022OptimizingLM]. Notable contributions include AGNN and EGNN  [ijcai2020p231, craandijk2022enforcement], which explores the use of graph neural network (GNN) to determine argumentation semantics, demonstrating that neural networks can effectively learn and predict the acceptability of arguments in various scenarios. These works highlight the promising intersection of neural networks and argumentation theory, however, the intrinsic black-box characteristic of neural networks leads to the lack of explainability in prediction. It is hard to show explicitly how their models predict the acceptability as their basis for decision-making, which is crucial in real applications.
AAF justifies argument acceptability through different semantics, thereby making selections among conflicting sets of arguments. Consequently, semantics computation is of great importance. In this paper, we focus on enhancing the conflict-solving capabilities of LLMs by leveraging abstract argumentation, integrating language model training with symbolic computation. Our objectives are twofold: first, to train an LLM capable of recognizing argumentation frameworks and assessing argument acceptability; second, to enhance the explainability of LLMs through this training, thereby potentially addressing some of their inherent limitations. To achieve these goals, we leverage the self-explanation capacity of LLMs [huang2023large, chuang2024faithlm] and train language models on a curated data corpus using supervised fine-tuning (SFT). Furthermore, we incorporate reinforcement learning from human feedback (RLHF) to further improve the interpretability and pedagogical quality of the model’s explanations [ouyang2022training, bakker2022fine, wong2024aligning]. This approach enables LLM to interpret its behavior during the problem-solving process through text generation. Specifically, we establish an AAF benchmark comprising formal frameworks that represent various conflict types. On the benchmark, we explore two fundamental tasks: computing the grounded extensions and complete extensions based on labelling algorithms. In the dataset, each sample comprises an AAF described using a graph description language, with the execution information of a recursive algorithm and justification process attached as the explanation. We then fine-tune two foundational models, Llama-3 [llama3] and Qwen2.5 [qwen25], on our established dataset by SFT and RLHF. The data organization embeds a post-explainability mechanism into the training process, which can reflect the LLM’s capability in understanding symbolic computation.
To summarize, our contributions are as follows:
  • 1.
    We establish an argumentation benchmark with algorithm explanation, enabling the training of explainable LLMs to argument acceptability using labeling algorithms.
  • 2.
    A methodology for integrating argumentation frameworks with language models is proposed through fine-tuning, improving the model’s ability to solve conflict and explain underlying computation dynamics.
  • 3.
2 Related Work
Within the domain of formal logical deduction using language models, the primary focus has predominantly been on first-order logic [tian2021diagnosing]. Works such as RuleTaker [Clark2020Transformers] and FOLIO [han2022folio] exemplify the application of artificial intelligence in deducing logical conclusions from structured premises, thereby demonstrating models’ capacity for formal logical reasoning. However, argumentative reasoning, which involves both logical reasoning and graph comprehension, has been largely overlooked [guo2023gpt4graph].
Previous research has attempted to integrate argumentation frameworks or with neural networks for interpretation [potyka2021interpreting] and semantics computation [craandijk2022enforcement, ijcai2020p231]. For instance, the work on interpreting neural networks as quantitative argumentation frameworks [potyka2021interpreting] explores how neural networks can be understood through the lens of argumentation. Similarly, the AGNN [ijcai2020p231] investigates the application of deep learning techniques to abstract argumentation semantics. Additionally, research in [craandijk2022enforcement] proposes enforcement heuristics for argumentation with deep reinforcement learning, aiming to enhance the decision-making process in argumentation frameworks. In [mileva2023unifying], researchers aim to consolidate diverse approaches to learning argumentation semantics within a unified framework.
Recent research has integrated argumentation techniques with large language models to enhance reasoning capabilities in various tasks. Görur et al. [gorur2024large] demonstrate that LLMs like Llama-2 and Mistral outperform traditional models in Relation-Based Argument Mining by effectively identifying support and attack relations between arguments. Chen et al. [chen2024exploring] conduct a comprehensive investigation into the potential of LLMs, such as ChatGPT, for computational argumentation. Their study focuses on argument mining and generation tasks, evaluating these models in both zero-shot and few-shot learning scenarios across diverse datasets. Freedman et al. [freedman2024argumentative] propose an innovative approach to augment the reasoning abilities of LLMs through the incorporation of argumentative reasoning, which leverages LLMs to construct argumentation frameworks. The integration of these methodologies holds significant promise for advancing the development of robust artificial intelligence systems capable of sophisticated argumentation and reasoning.
3 Preliminaries
@@
Abstract argumentation frameworks can be concisely represented by directed graphs, where arguments are nodes and edges represent the attack relation. We use the AAF in Figure 1 (denoted as Example 1) as our illustration example in this paper. The status of a given argument is determined through argumentation semantics, often yielding results in the form of extensions [dung1995acceptability] or labellings [wu2010labelling], which can be used interchangeably. In this work, we primarily use labellings, and the explanation of computation also involves concepts such as conflict-freeness.
@@
Definition 2
A labelling $\mathcal{L}$ for an abstract argumentation framework $\mathcal{F}=(\mathcal{A},\mathcal{R})$ is a function $\mathcal{L}:\mathcal{A}\rightarrow$ $\mathtt{IN}$, $\mathtt{OUT}$, $\mathtt{UNDEC}$}.
@@ 4 Dataset Construction
This work focuses on leveraging semantics computation with LLMs for conflict resolution, rather than improving argumentation computation methods. To construct a benchmark with algorithm explanations, we first generate randomly a variety of challenging argumentation frameworks following the methodology described in [ijcai2020p231]. In generation process, duplicate samples are removed and the ground truth semantics are computed. The number of arguments in an AAF ranges from 6 to 25. To encompass various levels of difficulty, we generate 3,000 samples for the training set and 100 samples for the test set for each number of arguments $n$, resulting in a total of 60,000 training samples and 2,000 test samples  [^1] https://github.com/dxbdxx/llm_af_computation. The grounded, preferred, stable, and complete semantics have an average of 1.0, 1.4, 1.1, and 2.4 extensions per AAF, respectively, with 4.3, 6.4, 7.6, and 6.0 arguments per extension, respectively.
@@ 4.1 Instruction and Problem
The instruction includes a basic introduction of AAF, the task description, and return format requirements. Since we are conducting formal reasoning, we ask LLMs to generate a JSON object at the end for clarity and concreteness. To render the abstract representation readable for both humans and LLMs, we adopt graph description languages to represent the original argumentation frameworks. We use three graph description languages: GraphML [brandes2002graphml], Graphviz DOT, and JSON. For each data sample, one format is randomly chosen during the problem statement generation.
@@ 4.2.1 Grounded Semantics
The labelling algorithm for generating the grounded labelling begins by assigning $\mathtt{IN}$ to all arguments that are not attacked. It then proceeds iteratively: $\mathtt{OUT}$ is assigned to any argument attacked by an argument that has just been made $\mathtt{IN}$, and $\mathtt{IN}$ is assigned to arguments whose attackers are all $\mathtt{OUT}$. The iteration continues until no new arguments can be assigned $\mathtt{IN}$ or $\mathtt{OUT}$. Any arguments that remain unlabelled are then assigned $\mathtt{UNDEC}$. This algorithm is sound and complete as it effectively mimics the construction of the grounded extension through the iteration of a framework’s characteristic function [modgil2009proof].
@@ 4.4 Enhancing Data Diversity
To further increase data diversity and reduce overfitting, we utilize GPT-4o [gpt4] to paraphrase and polish the natural language descriptions in both training and testing samples. For each instance, the instruction and explanatory texts are automatically rephrased into multiple diverse versions, while the formal graph representation (such as DOT, GraphML, or JSON format) and the answer remain unchanged.
5 Training and Baseline Methods
In this section, we introduce the training strategies and baseline methods used for argumentation semantics computation with LLMs.
5.1 Supervised Fine-Tuning
We mainly utilize the constructed benchmark dataset to perform SFT of LLMs for semantics computation. Each data instance follows an instruction-tuning format, where the instruction and problem fields are used as input, and the explanation (optional) and answer fields as supervised output. Following standard instruction tuning practices, we adopt a sequence-to-sequence training paradigm, enabling the model to generate both step-by-step reasoning and the final solution in a structured JSON format.
5.2 Chain-of-Thought Baseline
Prompting generally results in models making predictions with lower accuracy compared to models that have been fine-tuned with substantial training data [zhang2023large]. To validate this observation and establish a baseline method, we design a Chain-of-Thought (CoT) pipeline [NEURIPS2022_9d560961] that computes extensions based on a labelling algorithm. The argumentation computation is divided into multiple steps, framed as a reasoning task.
For computing grounded semantics, LLMs are taught to identify arguments and their relations, starting with those that aren’t attacked. They follow steps to label arguments until no more changes are needed. As for complete Semantics, the models take arguments that are labelled $\mathtt{UNDEC}$ and decide which can be further add to set $\mathtt{IN}$ to form a complete extension, checking if their choices make sense. The concrete prompts and implementation steps can be found in Appendix.
5.3 Integrating Human Feedback
The semantics computation based on labelling algorithms is primarily designed from a programmatic perspective. In contrast, when humans manually compute extensions, they often adopt a more intuitive and natural approach to explanation. Human explanations tend to be expressed in a way that is easier to understand. However, models trained solely via SFT typically generate outputs that closely follow the templates used during training, resulting in less flexible explanations.
To investigate the impact of human annotation on explanation quality, we design an experiment that incorporates reinforcement learning. Specifically, we provide both the problem statements and reference answers to expert annotators, asking them to play the role of a teacher and annotate step-by-step explanations as if instructing a student on how to compute extensions. The goal is to collect high-quality, pedagogical explanations that could potentially improve the explainability of model outputs. In this paper, we annotate grounded semantics and a portion of the data. More implementation details and annotated examples can be found in the Appendix.
6 Experiment
6.1 Implementation Details
In our training process, we employ the LoRA technique [hu2022lora] to facilitate efficient fine-tuning. Throughout all experiments, we utilize the bfloat16 (bf16) computation datatype. For model training, the maximum context length is set to 2048 tokens. The learning rate is set to $1\text{e}{-}4$ for SFT, and the training time is around 10 hours. For RLHF, the learning rate is $1\text{e}{-}6$. The LLMs are trained on a server with eight NVIDIA A100-SXM-40GB GPUs. For inference, we use vLLM [kwon2023efficient] and set the temperature to 0.3. It takes 0.4 seconds to generate all responses for semantics computation on a single GPU.
@@ 6.1 Implementation Details
Qwen2.5-SFT 0.97 0.99 0.91 0.99 0.96 0.88 0.98 0.96 0.90 0.97 0.95
Llama3-SFT 0.95 0.99 0.85 0.99 0.89 0.86 0.90 0.89 0.88 0.88 0.89
@@
Figure 4: Learning curves of semantics prediction.
6.2 Metrics
To evaluate the performance of the models, we consider two perspectives: the extension view and the argument view.
Extension View Accuracy
From the extension view, we define the accuracy of extension prediction, denoted as ACCσ. This metric represents the proportion of cases where the predicted $\mathtt{IN}(\mathcal{L})$ exactly matches the ground truth extension for a given semantics $\sigma$. We also use the Pass@k metric to evaluate model robustness in predicting extensions.
Argument View Accuracy
From the argument view, we assess the acceptability of arguments by calculating their credulous and skeptical acceptability based on the predicted labellings. For argument acceptability classification, we use the Matthews Correlation Coefficient (MCC), which is particularly suited for imbalanced datasets. The metrics are denoted as MCCc for credulous acceptability and MCCs for skeptical acceptability, respectively.
To evaluate the effectiveness of conflict resolution, we employ the following two metrics:
Conflict-Free Proportion
This metric measures the average proportion of generated extensions that satisfy the conflict-freeness property across all argumentation frameworks in the dataset. The conflict-freeness property ensures that no two arguments in a single extension attack each other. Formally, for a dataset with $N$ argumentation frameworks, the Conflict-Free Proportion (CFP) is defined as:
$\text{CFP}=\frac{1}{N}\sum_{\mathcal{E}}\frac{|\{E\in\mathcal{E}\mid E\text{ is conflict-free}\}|}{|\mathcal{E}|}$
where $\mathcal{E}$ represents the set of generated extensions.
At Least One Successful Extension
The metric evaluates whether at least one of the generated extensions is valid. For a given semantics $\sigma$ and a set of generated extensions $\mathcal{E}$, the At Least One Successful Extension (ALSE) for a single argumentation framework $\mathcal{F}$ is defined as:
$\text{ALSE}_{\mathcal{F}}(\mathcal{E})=\begin{cases}1,&\text{if }\exists E\in\mathcal{E}_{\mathcal{F}}\text{ such that }E\in\sigma(\mathcal{F}),\\ 0,&\text{otherwise}.\end{cases}$
The overall ALSE score for a dataset is then computed as the proportion of cases where at least one valid extension is generated:
$\text{ALSE}=\frac{1}{N}\sum_{\mathcal{F}}\text{ALSE}_{\mathcal{F}}(\mathcal{E}_{\mathcal{F}})$
where $\mathcal{E}_{\mathcal{F}}$ represents the set of generated extensions for the framework $\mathcal{F}$.
@@ 6.3 Main Results
It is worth noting that the accuracy of our models in prf prediction is comparable to, but does not exceed, that of GNN-based approaches [ijcai2020p231], which benefit from architectures well-suited to argumentation tasks. However, a key advantage of our method lies in its ability to directly compute extensions and provide corresponding explanations, offering a comprehensive understanding of the argumentation process. This capability underscores the practical applicability of our model in diverse semantic contexts.
@@
6.4 Effectiveness on Conflict Resolving
Models com prf
CFP ALSE CFP ALSE
Qwen2.5-7B 0.71 0.62 0.78 0.63
Llama3-8B 0.51 0.42 0.45 0.48
Qwen2.5-SFT 1.00 0.99 1.00 0.95
Llama3-SFT 1.00 0.99 1.00 0.86
Table 3: The conflict resolving ability comparison of LLMs.
The evaluation results in Table 3 notably demonstrate the differences in conflict resolution capabilities before and after training. Before training, Qwen2.5B achieves CFP scores of 0.71 and 0.78, and ALSE scores of 0.62 and 0.63 under the complete and preferred semantics, respectively. In comparison, Llama3 shows lower performance, with CFP scores of 0.51 and 0.45, and ALSE scores of 0.42 and 0.48 in the same scenarios. After training, both models achieve near-perfect results: Qwen2.5-SFT attains CFP and ALSE scores of 1.00 and 0.99 under com, and 1.00 and 0.95 under prf; while Llama3-SFT reaches 1.00 and 0.99 under com, and 1.00 and 0.86 under prf. These results clearly demonstrate that training significantly enhances the models’ abilities to generate logically consistent and semantically valid solutions in argumentation frameworks.
6.5 Robustness
Figure 5: Robustness: Pass@k vs. Noisy Data Ratio
We generate our dataset through symbolic computation to ensure clean data. To evaluate the robustness of our trained models against noise, we manually create corrupted datasets by generating erroneous samples for grounded semantics. Specifically, to generate an error case, we randomly select an intermediate set in a clean derivation and replace half of its elements with random values. For example, we may modify the label set IN from $\{1,3,6,7\}$ to $\{2,3,5,7\}$, and then recompute OUT and UNDEC accordingly. Such a step error can propagate to the final answer, thereby introducing noise into the reasoning process.
To assess the stability of model outputs, we train the models on the corrupted datasets and report Pass@k performance in Figure 5 as the proportion of noisy samples increases. As shown in the results, both models exhibit a gradual decrease in Pass@k as the proportion of noisy samples increases, and Qwen2.5 consistently outperforms Llama3 under all noise levels.
6.6 Direct vs. Indirect Complete Semantics Computation
Figure 6: Accuracy of two-step vs. end-to-end prediction of complete extensions.
In the task of predicting the complete extensions, we divide the computation into two steps: first computing the grounded extension, and then computing the complete extension. In the second step, we provide the grounded extension as a known condition in the prompt. If we adopt an end-to-end approach to predict the complete extensions, errors can accumulate across steps. In real-world applications, however, such end-to-end prediction is often more meaningful. Figure 6 shows the results for both methods. For each model, the left bar represents the two-step approach, while the right bar represents the end-to-end approach. The annotated gap $\Delta$ indicates the accuracy drop caused by error accumulation in the end-to-end setting.
6.7 Effect of Graph Input
AAF structures can be visualized as directed graphs for model input (see Figure 1). To investigate the effect of visualized graph inputs, we conduct experiments using Qwen2.5-VL-7B [qwen25vl], a multi-modal model comparable in scale to Qwen2.5-7B. For comparison, we evaluate both unimodal (text-only) and multimodal (visual+text) settings. In the Visual+Text setting, the problem description is converted into an image, while other components remain unchanged. Table 4 summarizes the performance under different input formats. For the Qwen2.5-VL-SFT model, using visualized graph inputs together with text achieves slightly higher accuracy on both grounded extension and complete extension prediction compared to the text-only setting. This suggests that incorporating visual representations of AAF structures can provide additional benefits in understanding and reasoning, although the improvement is marginal in this case.
6.8 Errors in Computing Grounded Semantics
In this part, we summarize the errors and failures encountered by the LLMs when computing the grounded labelling.
Model Input Format $_{\textsf{grd}} ACC$$_{\textsf{com}}$
Qwen2.5-7B Text 0.31 0.43
Qwen2.5-SFT Text 0.97 0.85
Qwen2.5-VL-SFT Text 0.98 0.87
Qwen2.5-VL-SFT Visual+Text 0.98 0.88
Table 4: Performance comparison under different input formats.